# HierarchicalPrune: Position-Aware Compression for Large-Scale Diffusion Models

### v1

@@ 3.2 Main Results
Image Quality: Fig. 6 presents a qualitative comparison of example outputs. HierarchicalPrune shows better visual outputs, outperforming all the baselines, and maintaining high fidelity to the original DMs, both in structure and fine details (see more examples in Fig. 2 and Appendix C).
Table 1 shows that HierarchicalPrune achieves comparable quality to the original model based on both HPSv2 and GenEval. While baselines induce substantial quality degradation of 38.5-41.2%, our approach yields only a minimal drop ($\sim$3.2-4.8%), showcasing the effectiveness of our hierarchical compression strategy.
Table 1: Quantitative evaluation of image quality. KOALA and BK-SDM experience substantial degradation of image quality when reducing memory usage by 20-30%. HierarchicalPrune achieves significantly reduced memory usage while maintaining the image quality close to the original models.
@@
Linear DiT SANA-Sprint (100%) 0.77 29.61 -
Original (100%) 0.71 30.29 -
KOALA (79.4%) 0.37 19.99 38.5%
SD3.5 BK-SDM (79.4%) 0.38 21.21 41.2%
Large Turbo Ours (HPP, PWP, Quant) (22.5%) 0.69 28.15 4.8%
Ours (All) (20.5%) 0.62 26.29 13.3%
Original (100%) 0.66 29.71 -
FLUX.1 KOALA (70.5%) 0.56 24.04 17.1%
Schnell BK-SDM (70.5%) 0.57 24.67 15.3%
Ours (All) (19.6%) 0.64 28.69 3.2%
@@
Table 2: Comparison of the peak memory during inference and per-step latency of the DM Transformer measured on an A6000 GPU with 48 GB of VRAM.
Model Method Memory $\Downarrow$ Reduction $\Uparrow$ Latency $\Downarrow$ Reduction $\Uparrow$
Linear DiT SANA-Sprint 3.14 GB - 54ms -
Original 15.8 GB - 823 ms -
SD3.5 KOALA 12.6 GB 20.6% 642 ms 22.0%
Large Turbo BK-SDM 12.6 GB 20.6% 642 ms 22.0%
Ours (All) 3.24 GB 79.5% 593 ms 27.9%
Original 22.6 GB - 756 ms -
FLUX.1 KOALA 15.9 GB 29.5% 432 ms 42.9%
Schnell BK-SDM 15.9 GB 29.5% 432 ms 42.9%
Ours (All) 4.44 GB 80.4% 469 ms 38.0%
Peak Memory & Latency: On the resource front, as in Table 2, our compression pipeline achieves 79.5% lower peak memory (from 15.8 GB to just 3.24 GB) and 27.9% faster latency compared to the original SD3.5 Large Turbo model. Measurements of memory and latency were conducted on both server- and desktop-grade GPUs with different specifications (A100, A6000, and GTX 3090 with 80, 48, and 25 GB VRAM, respectively, reported in Appendix Table 5). While we report the measurement results from A6000, the reduction rates are similar across GPUs. Notably, the peak memory of our compressed SD3.5 model (3.24 GB) is comparable to SANA-Sprint (3.14 GB), but with substantially better image quality. Similarly, for FLUX.1-Schnell, HierarchicalPrune achieves 80.4% peak memory reduction and 38.0% speed gain, indicating its generalisability to larger DMs.
@@ 3.3 Ablation Study and Analysis
Impact of Each Component of HierarchicalPrune: We conducted an ablation study of HierarchicalPrune to investigate the contribution of each component: (1) HPP, (2) PWP, (3) HPP+PWP+SGDistill, and (4) our final form, HPP+PWP+SGDistill+Quant. Table 3 shows that using HPP+PWP drastically improves image quality compared to HPP only and prior works such as KOALA and BK-SDM. Moreover, SGDistill substantially prevents the image generation quality degradation compared to HPP+PWP with an aggressive paramter reduction rate of 30%, leading to higher memory reduction without compromising performance drastically. Our method - leveraging all three components - outperforms all the baselines with superior image quality, lower memory, and faster execution.
@@
Method Memory (%) GenEval $\Uparrow$ HPSv2 $\Uparrow$ Avg. Reduction $\Downarrow$
Original (100%) 0.71 30.29 -
Ours (HPP) (79.4%) 0.03 11.08 79.4%
Ours (+PWP) (79.4%) 0.71 29.97 2.5%
Ours (+Quant) (22.5%) 0.69 28.15 4.8%
Ours (HPP) (71.5%) 0.0 7.00 88.4%
Ours (+PWP) (71.5%) 0.46 21.74 31.9%
Ours (+SGDistill) (71.5%) 0.64 27.29 10.1%
Ours (+Quant) (20.5%) 0.62 26.29 13.3%
@@ 4 Related Work
T2I Diffusion Models. Since Stable Diffusion (SD) (Rombach et al. 2022), there has been rapid community adoption and iterative updates in the field. SDv1.4 and v1.5 were released in 2022, enhancing efficiency and enabling specialised fine-tuning (e.g., DreamBooth). By mid-2023, Stable Diffusion XL (SDXL) significantly advanced resolution ($\times1024$), text comprehension, and image quality. By the end of 2024, the adoption of MMDiT (Esser et al. 2024) as a backbone in SD3, SD3.5, and FLUX brought a significant boost in image generation quality and alignment with long text input. At the same time, the increasing parameter count, particularly in the backbone Transformer blocks, significantly improved image quality but with excessive resource demands. As SOTA models like SD3.5 scale up to as many as 8B parameters, efficient inference on resource-constrained devices becomes impractical. This underlines the significance of approaches such as our HierarchicalPrune framework, to enable DM deployment outside of high-end compute setups.
Compression Techniques. Traditional model compression methods, including distillation (Hinton, Vinyals, and Dean 2015), pruning (Han, Mao, and Dally 2016) (such as depth pruning via block removal (Ghiasi, Lin, and Le 2018; Kim et al. 2024b), width pruning (Kwon et al. 2024)), and quantisation (Jacob et al. 2018) have been applied to DMs (Lee et al. 2024; Kim et al. 2024a; Castells et al. 2024b; Hu et al. 2024a; Fang et al. 2024; Li et al. 2025), resulting in variable parameter efficiency gains. Specifically, (Kim et al. 2024a) and (Lee et al. 2024) explore block removal in the U-Net backbone of SD1.5 and SDXL models, respectively. More recently, (Li et al. 2025) explored the effectiveness of quantisation on DMs for memory reduction, hence our HierarchicalPrune employs an adjunct post-training quantisation feature as part of its design. In this paper, we investigated the hierarchical nature of the SOTA MMDiT backbone in SD3.5 and FLUX, and further leveraged these insights for the compression of SOTA multi-billion-scale DMs, under a unified framework.
@@ A.1 Datasets
We use the YE-POP dataset from HuggingFace for distillation training of the target models, similar to prior work by (Lee et al. 2024). The YE-POP dataset is comprised of 500,000 high-quality images sampled from the LAION-5B dataset (Schuhmann et al. 2022), incorporating a variety of visual modalities across multiple disciplines, including images of natural scenes, objects, people, and artistic images, providing extensive breadth for the training and evaluation of diffusion models (DMs). Our dataset preprocessing pipeline includes standard normalisation processes and low-resolution image and damage filtering, and ensures consistent image resolutions and aspect ratios to provide a fair data comparison for our model configurations and baseline methods.
@@ A.2 Architectures and Implementation
Implementation: The HierarchicalPrune pipeline is implemented in PyTorch. In particular, we built on top of checkpoints and inference pipelines for SD3.5 Large Turbo and FLUX.1-Schnell from the Diffusers library  (von Platen et al. 2022). This allows us both consistency with pre- and post-processing steps, and also reproducible experiments. The distillation training process is conducted on a single node of 8 A100 GPUs with 80 GB of VRAM.
Quantization Strategy: To further reduce the memory footprint of the models, we employed 4-bit weight quantization from using bitsandbytes (Dettmers et al. 2022, 2023).
@@ A.3 Baseline Comparisons
BK-SDM (Kim et al. 2024a): BK-SDM performs block pruning for U-Net-based diffusion using CLIP (Hessel et al. 2021) scores as importance metric, where it ranks blocks based on the reduction of CLIP scores of the output images from the model before and after pruning each block. After pruning, BK-SDM distils the compressed model using knowledge from the original unpruned model, to recover performance from the previous steps.
KOALA (Lee et al. 2024): KOALA uses an importance ranking approach by calculating input-output cosine similarity to inform the block pruning strategy for U-Net based DMs. The blocks of the transformer have intrinsic importance relative to input-output transformations by evaluating the cosine similarity between input and output representations/features rather than utilising an external quality metric of transformation.
SANA-Sprint-1.6B (Chen et al. 2025b): We use SANA-Sprint-1.6B as the state-of-the-art (SOTA) baseline for an efficient, small-scale diffusion purpose-built model, incorporating techniques such as deep-compression auto-encoder (Chen et al. 2025a) and linear attention (Katharopoulos et al. 2020). Intrinsically, SANA models require training of the transformer from scratch, which departs from compressing existing models.
@@ A.4 Evaluation Framework
Quantitative Metrics: We adopt two representative and complementary image quality assessment metrics, capturing two different aspects of the quality of generation. First, GenEval (Ghosh, Hajishirzi, and Schmidt 2023) provides an all-encompassing evaluation across three areas, i.e., semantic fidelity, visual fidelity, and prompt adherence. GenEval is now widely used for evaluations of T2I models (Xie et al. 2025; Esser et al. 2024). Second, HPSv2 (Wu et al. 2023a) provides a contrasting perspective on image quality assessment, with demonstrated effectiveness in estimating the perceptual preferences of humans about image quality. HPSv2 has been utilised in several recent works such as KOALA (Lee et al. 2024).
Human Preference Metrics: We conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics (see Appendix A.5 for further details regarding our user study design). Similarly to (Sheynin et al. 2023) and (Dai et al. 2023), we evaluate two aspects: (1) Text Alignment: How well a generated image follows the description of the text prompt, (2) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors.
@@ A.5 Detailed User Study Design
Acknowledging the inconsistency of automated metrics, especially in discerning fine aspects of image quality and preferences for aesthetics, we choose to conduct a full user study of methods following the protocol of (Sheynin et al. 2023) and (Dai et al. 2023). In our user evaluation method, we are focused on measuring two aspects of user experience:
@@
Figure 8: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024). (b,c,d,e,f) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent.
@@ Appendix B Additional Analysis
Figure 10: Fine-grained contribution analysis of FLUX.1-Schnell on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024). (b,c,d,e,f,g,h) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent. Note that the first part of the model from block 0 to 18 is composed of the transformer blocks with five components (b-f) and the second part of the model is made of single transformer blocks with three components (b,g,h).
@@ C.1 Computational Overhead Analysis
Overall, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling. While substantial, this represents a one-time cost that is amortised across all deployments. Furthermore, training small-scale DMs (1-2B parameter-sized) from scratch requires 140,000 and/or 200,000 A100 GPU hours for smaller models such as SD1.4 and SD2.1, respectively, as reported in (Pernias et al. 2024), let alone large-scale DMs (8-11B parameter-sized), which could require significantly more GPU hours to train from scratch. Yet, our contribution analysis cost equals <0.006% of typical DM training.
@@
KOALA (79.4%) 0.37 19.99 ($\pm$ 0.074) 38.5%
BK-SDM (79.4%) 0.38 21.21 ($\pm$ 0.077) 41.2%
Ours (HPP) (79.4%) 0.03 11.08 ($\pm$ 0.042) 79.4%
SD3.5 Ours (+PWP) (79.4%) 0.71 29.97 ($\pm$ 0.077) 2.5%
Large Turbo Ours (+Quant) (22.5%) 0.69 28.15 ($\pm$ 0.078) 4.8%
@@
FLUX.1 KOALA (70.5%) 0.56 24.04 ($\pm$ 0.075) 17.1%
Schnell BK-SDM (70.5%) 0.57 24.67 ($\pm$ 0.074) 15.3%
@@
Overall Computational Requirement: When considered in its entirety, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling across different model architectures, and the subsequent distillation pipeline demands 603-1260 A100 GPU hours. While substantial, these represent one-time costs that amortise across all subsequent deployments. To place this computational investment in perspective, training small-scale DMs (1-2B parameters) from scratch requires 140,000-200,000 A100 GPU hours for SD1.4 and SD2.1 models, respectively, as reported (Pernias et al. 2024), let alone large-scale DMs (8-11B parameters), which would require considerably more computational resources. Notably, our contribution analysis cost equals less than 0.006% of typical DM training overhead, while our distillation training process represents less than 0.30-0.63% of the computational cost required to train comparable models from scratch.

### v2

@@ 3.2 Main Results
Image Quality under Memory Compression: Fig. 6 presents a qualitative comparison of example outputs. HierarchicalPrune shows better visual outputs, outperforming all the baselines, and maintaining high fidelity to the original DMs, both in structure and fine details. More examples can be found in Fig. 2 and Appendix C. Table 1 shows that HierarchicalPrune achieves comparable quality to the original model based on both HPSv2 and GenEval. While baselines induce substantial quality degradation of 38.2-45.1%, our approach yields only a minimal drop ($\sim$3.2-4.8%), given the same range of memory compression ratio, showcasing the effectiveness of our hierarchical compression strategy.
Importantly, HierarchicalPrune achieves 79.5% lower peak memory, a reduction from 15.8 GB to 3.24 GB for SD3.5 Large Turbo model. Notably, the peak memory of our compressed SD3.5 model (3.24 GB) is comparable to SANA-Sprint (3.14 GB), but with substantially better image quality. For FLUX.1-Schnell, our method achieves 80.4% peak memory reduction, indicating its generalisability to larger DMs.
Table 1: Image quality measured in GenEval and HPSv2 scores and the corresponding peak memory measurements in GB and remaining ratios (%) after compression. KOALA and BK-SDM experience substantial degradation of image quality when reducing memory usage by 20-30%. HierarchicalPrune achieves significantly reduced memory usage while maintaining the image quality close to the original models.
@@
Linear DiT SANA-Sprint 3.14 GB (100%) 0.77 29.61 -
Original 15.8 GB (100%) 0.71 30.29 -
KOALA 12.6 GB (79.4%) 0.37 19.99 41.2%
KOALA (+Quant) 3.56 GB (22.5%) 0.33 18.44 46.4%
SD3.5 BK-SDM 12.6 GB (79.4%) 0.38 21.21 38.2%
Large Turbo BK-SDM (+Quant) 3.56 GB (22.5%) 0.34 19.83 43.3%
Ours (HPP+PWP+Q) 3.56 GB (22.5%) 0.69 28.15 4.8%
Ours (All) 3.24 GB (20.5%) 0.62 26.29 13.3%
Original 22.6 GB (100%) 0.66 29.71 -
FLUX.1 KOALA 15.9 GB (70.5%) 0.38 25.24 28.7%
Schnell BK-SDM 15.9 GB (70.5%) 0.45 27.38 19.8%
Ours (All) 4.44 GB (19.6%) 0.64 28.69 3.2%
@@
Training Cost Comparison: Beyond quality preservation, HierarchicalPrune requires only 615–1,287 A100 GPU hours for architectural profiling and distillation (Appendix C.1), compared to 140k–200k A100 GPU hours required to pre-train small-scale DMs like SD1.4 and SD2.1 (2024ICLRgU58d5QeGv_cascade). This demonstrates that compressing existing large models is far more cost-efficient than training compact alternatives from scratch, while maintaining superior quality.
Table 2: Comparison of per-step latency of the DM Transformer measured on an A6000 GPU with 48 GB of VRAM.
Model Method Latency $\Downarrow$ Reduction $\Uparrow$
Linear DiT SANA-Sprint 54 ms -
Original 823 ms -
KOALA 642 ms 22.0%
SD3.5 BK-SDM 642 ms 22.0%
Large Turbo Ours (HPP, PWP, Quant) 593 ms 27.9%
Ours (All)
Original 756 ms -
FLUX.1 KOALA 432 ms 42.9%
Schnell BK-SDM 432 ms 42.9%
Ours (All) 469 ms 38.0%
Latency: We measure per-step latency of HierarchicalPrune and baselines both server- and desktop-grade GPUs with different specifications (A100, A6000, and GTX 3090 with 80, 48, and 25 GB VRAM, respectively, reported in Appendix Table 5). While we report the measurement results from A6000 in Table 2, the reduction rates are similar across GPUs. Our compression pipeline achieves 27.9% and 38.0% latency reduction compared to the original SD3.5 Large Turbo and FLUX.1-Schnell, respectively.
@@ 3.3 Ablation Study and Analysis
Impact of Each Component of HierarchicalPrune: We conducted an ablation study of HierarchicalPrune to investigate the contribution of each component: (1) HPP, (2) PWP, (3) HPP+PWP+SGDistill, and (4) our final form, HPP+PWP+SGDistill+Quant. Table 3 shows that using HPP+PWP drastically improves image quality compared to HPP only and prior works (i.e., BK-SDM, KOALA). Moreover, by leveraging intra-block sensitivity, SGDistill substantially reduces image quality degradation from 31.0% to 10.1% compared to HPP+PWP at an aggressive paramter reduction rate of 30% [^2] We set $r_{thres}=0.25$ for both models, determined where quality degradation exceeded 15% without SGDistill.. Our method outperforms all baselines with superior image quality, lower memory, and faster execution.
@@
Prunning
Ratio
Method
Remaining
Memory (%)
GenEval $\Uparrow$ HPSv2 $\Uparrow$ Reduction $\Downarrow$
None (0%) Original 100% 0.71 30.29 -
Moderate Pruning (20%) Ours (HPP) 79.4% 0.03 11.08 79.4%
Ours (+PWP) 79.4% 0.71 28.97 2.5%
Ours (+Quant) 22.5% 0.69 28.15 4.8%
Aggressive Pruning (30%) Ours (HPP) 71.5% 0.0 7.00 88.4%
Ours (+PWP) 71.5% 0.46 21.74 31.9%
Ours (+SGDistill) 71.5% 0.64 27.29 10.1%
Ours (+Quant) 20.5% 0.62 26.29 13.3%
@@ 4 Related Work
T2I Diffusion Models. Since Stable Diffusion (SD) (2022CVPRLDM), there has been rapid community adoption and iterative updates in the field. SDv1.4 and v1.5 were released in 2022, enhancing efficiency and enabling specialised fine-tuning (e.g., DreamBooth). By mid-2023, Stable Diffusion XL (SDXL) significantly advanced resolution ($\times1024$), text comprehension, and image quality. By the end of 2024, the adoption of MMDiT (esser2024mmdit) as a backbone in SD3, SD3.5, and FLUX brought a significant boost in image generation quality and alignment with long text input. At the same time, the increasing parameter count, particularly in the backbone Transformer blocks, significantly improved image quality but with excessive resource demands. As SOTA models like SD3.5 scale up to as many as 8B parameters, efficient inference on resource-constrained devices becomes impractical. This underlines the significance of approaches such as our HierarchicalPrune framework, to enable DM deployment outside of high-end compute setups.
Compression Techniques. Traditional model compression methods, including distillation (hinton2015distilling), pruning (han2015deepcompression) (such as depth pruning via block removal (ghiasi2018dropblock; kim2024layermerge), width pruning (kwon2024tinytrain)), and quantisation (jacob2018quantization) have been applied to DMs (lee2024koala; kim2024bksdm; castells2024edgefusion; hu2024snapgen; fang2024tinyfusion; li2025svdquant; kim2024IDcompression; zhang2024laptopdiff), resulting in variable parameter efficiency gains. Specifically, (kim2024bksdm) and (lee2024koala) explore block removal in the U-Net backbone of SD1.5 and SDXL models, respectively. More recently, (li2025svdquant) explored the effectiveness of quantisation on DMs for memory reduction, hence our HierarchicalPrune employs adjunct post-training quantisation as part of its design. Orthogonal to compression, $\Delta$-DiT (chen2024deltaDit) identifies hierarchical patterns in DiT (MMDiT predecessor) for inference caching. Our analysis confirms that this generalises to modern MMDiT and further leverages these insights for the compression of SOTA multi-billion-scale DMs, under a unified framework.
@@ A.1 Datasets
We use the YE-POP dataset from HuggingFace for distillation training of the target models, similar to prior work by (lee2024koala). The YE-POP dataset is comprised of 500,000 high-quality images sampled from the LAION-5B dataset (2022NeurIPSLAION5B), incorporating a variety of visual modalities across multiple disciplines, including images of natural scenes, objects, people, and artistic images, providing extensive breadth for the training and evaluation of diffusion models (DMs). Our dataset preprocessing pipeline includes standard normalisation processes and low-resolution image and damage filtering, and ensures consistent image resolutions and aspect ratios to provide a fair data comparison for our model configurations and baseline methods.
@@ A.2 Architectures and Implementation
Implementation: The HierarchicalPrune pipeline is implemented in PyTorch. In particular, we built on top of checkpoints and inference pipelines for SD3.5 Large Turbo and FLUX.1-Schnell from the Diffusers library  (von-platen-etal-2022-diffusers). This allows us both consistency with pre- and post-processing steps, and also reproducible experiments. The distillation training process is conducted on a single node of 8 A100 GPUs with 80 GB of VRAM.
Quantization Strategy: To further reduce the memory footprint of the models, we employed 4-bit weight quantization from using bitsandbytes (dettmers2022llmint8; dettmers2023qlora).
@@ A.3 Baseline Comparisons
BK-SDM (kim2024bksdm): BK-SDM performs block pruning for U-Net-based diffusion using CLIP (hessel-etal-2021-clipscore) scores as importance metric, where it ranks blocks based on the reduction of CLIP scores of the output images from the model before and after pruning each block. After pruning, BK-SDM distils the compressed model using knowledge from the original unpruned model, to recover performance from the previous steps.
KOALA (lee2024koala): KOALA uses an importance ranking approach by calculating input-output cosine similarity to inform the block pruning strategy for U-Net based DMs. The blocks of the transformer have intrinsic importance relative to input-output transformations by evaluating the cosine similarity between input and output representations/features rather than utilising an external quality metric of transformation.
SANA-Sprint-1.6B (chen2025sanasprintonestepdiffusioncontinuoustime): We use SANA-Sprint-1.6B as the state-of-the-art (SOTA) baseline for an efficient, small-scale diffusion purpose-built model, incorporating techniques such as deep-compression auto-encoder (chen2025deep) and linear attention (linear_attn_katharopoulos2020). Intrinsically, SANA models require training of the transformer from scratch, which departs from compressing existing models.
@@ A.4 Evaluation Framework
Quantitative Metrics: We adopt two representative and complementary image quality assessment metrics, capturing two different aspects of the quality of generation. First, GenEval (ghosh2023geneval) provides an all-encompassing evaluation across three areas, i.e., semantic fidelity, visual fidelity, and prompt adherence. GenEval is now widely used for evaluations of T2I models (xie2025sana; esser2024mmdit). Second, HPSv2 (wu2023human) provides a contrasting perspective on image quality assessment, with demonstrated effectiveness in estimating the perceptual preferences of humans about image quality. HPSv2 has been utilised in several recent works such as KOALA (lee2024koala).
Human Preference Metrics: We conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics (see Appendix A.5 for further details regarding our user study design). Similarly to (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration), we evaluate two aspects: (1) Text Alignment: How well a generated image follows the description of the text prompt, (2) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors.
@@ A.5 Detailed User Study Design
Acknowledging the inconsistency of automated metrics, especially in discerning fine aspects of image quality and preferences for aesthetics, we choose to conduct a full user study of methods following the protocol of (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration). In our user evaluation method, we are focused on measuring two aspects of user experience:
@@
Figure 8: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion). (b,c,d,e,f) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent.
@@ Appendix B Additional Analysis
Figure 10: Fine-grained contribution analysis of FLUX.1-Schnell on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion). (b,c,d,e,f,g,h) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent. Note that the first part of the model from block 0 to 18 is composed of the transformer blocks with five components (b-f) and the second part of the model is made of single transformer blocks with three components (b,g,h).
@@ C.1 Computational Overhead Analysis
Overall, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling. While substantial, this represents a one-time cost that is amortised across all deployments. Furthermore, training small-scale DMs (1-2B parameter-sized) from scratch requires 140,000 and/or 200,000 A100 GPU hours for smaller models such as SD1.4 and SD2.1, respectively, as reported in (2024ICLRgU58d5QeGv_cascade), let alone large-scale DMs (8-11B parameter-sized), which could require significantly more GPU hours to train from scratch. Yet, our contribution analysis cost equals <0.006% of typical DM training.
@@
KOALA (79.4%) 0.37 19.99 ($\pm$ 0.074) 41.2%
KOALA (+Quant) (22.5%) 0.33 18.44 ($\pm$ 0.075) 46.4%
BK-SDM (79.4%) 0.38 21.21 ($\pm$ 0.077) 38.2%
BK-SDM (+Quant) (22.5%) 0.34 19.83 ($\pm$ 0.079) 43.3%
SD3.5 Ours (HPP) (79.4%) 0.03 11.08 ($\pm$ 0.042) 79.4%
Large Turbo Ours (+PWP) (79.4%) 0.71 28.97 ($\pm$ 0.077) 2.5%
Ours (+Quant) (22.5%) 0.69 28.15 ($\pm$ 0.078) 4.8%
@@
FLUX.1 KOALA (70.5%) 0.38 25.24 ($\pm$ 0.075) 28.7%
Schnell BK-SDM (70.5%) 0.45 27.38 ($\pm$ 0.074) 19.8%
@@
Overall Computational Requirement: When considered in its entirety, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling across different model architectures, and the subsequent distillation pipeline demands 603-1260 A100 GPU hours. While substantial, these represent one-time costs that amortise across all subsequent deployments. To place this computational investment in perspective, training small-scale DMs (1-2B parameters) from scratch requires 140,000-200,000 A100 GPU hours for SD1.4 and SD2.1 models, respectively, as reported (2024ICLRgU58d5QeGv_cascade), let alone large-scale DMs (8-11B parameters), which would require considerably more computational resources. Notably, our contribution analysis cost equals less than 0.006% of typical DM training overhead, while our distillation training process represents less than 0.30-0.63% of the computational cost required to train comparable models from scratch.